OpenStudy (ggrree):

Show that e^(-2ln5) = -25

OpenStudy (ggrree):

I understand that -2ln(5) = ln (5^-2), but I'm stuck from here.

OpenStudy (ggrree):

sorry, I mistyped the question. "Show that e^(-2ln5) = 1/25" not -25.

OpenStudy (anonymous):

e^ ln (5^-2) = 5^-2 ( by definition)

OpenStudy (anonymous):

5^-2 = 1/25

OpenStudy (ggrree):

Why is e^ ln (5^-2) = 5^-2 ?

OpenStudy (anonymous):

because ln = log to the base e

OpenStudy (anonymous):

it might help to consider logs to the base 10 log _10 100 = 2 because 10^2 = 100 therefore, substituting 2 = log _10 100 in 10^2 = 100: 10 ^ (log_10 100) = 100